OpenStudy (baldymcgee6):

How do I find the derivative of: f (t) = 125e^((1/6)*ln(38/25)*t)

OpenStudy (anonymous):

\[f(t) = 125e ^{1/6} \times \ln |(38/25)t|\]

OpenStudy (anonymous):

Use your product rule to split the function into 2 parts 1)\[f(t)=125e^{1/6}\] 2)\[g(t)=\ln(38/25)t\] \[f \prime (t)=125/6 \times e^{1/6}\] \[g \prime (t)=\ln(38/25)+1/t\] Now work the product rule. By the way, I used log rules to split g(t), then took the derivative.